OpenStudy (anonymous):

if 1 us dollar = .678 euros what is 44 euros worth in us dollars?

OpenStudy (anonymous):

x= 1 dollar x 44 euros/ 0.78 euros

OpenStudy (anonymous):

x= 44/0.678 = $64.89

hero (hero):

\[\frac{\text{1 us dollars}}{\text{x us dollars}} = \frac{\text{.678 euros}}{\text{44 euros}}\] \[\frac{\text{1}}{\text{x}} = \frac{\text{.678}}{\text{44}}\] \[.678x = 44\] \[x = \frac{44}{.678}\]

hero (hero):

\[\approx $64.90 \]

OpenStudy (mathstudent55):

The conversion: \( 1~ US$ = 0.678 ~euro\) Can be written as these two conversion fractions that equal 1. \(\dfrac{1~US$}{0.678~ euro} = \dfrac{0.678 ~euro}{1~US$} = 1\) Since the conversion fractions are equal to one, you can multiply any amount by them without changing the amount. You are only changing the form the amount is written in. Use the conversion fraction that will eliminate the unit you don't want: \(44 ~euro \times \dfrac{1~US$}{0.678~ euro} \) As you can see, the euro units will cancel out, and the answer will be in US $.

OpenStudy (anonymous):

@chad i think we can do your problem